OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, $135. This year, he paid $189. What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

woah

OpenStudy (kropot72):

Increase in ticket price = 189 - 135 = $54 Percentage increase is:\[\frac{54}{135}\times100=40 \]percent